# Harnessing the Universal Geometry of Embeddings

### v3

@@ Abstract
Figure 1: Left: input embeddings from different model families (T5-based GTR [46] and BERT-based GTE [32]) are fundamentally incomparable. Right: given unpaired embedding samples from different models on different texts, our model learns a latent representation where they are closely aligned.
@@ 1 Introduction
The Platonic Representation Hypothesis [18] conjectures that all image models of sufficient size converge to the same latent representation. We propose a stronger, constructive version of this hypothesis for text models: the universal latent structure of text representations can be learned and, furthermore, harnessed to translate representations from one space to another without any paired data or encoders.
@@
We draw inspiration from research on aligning word embeddings across languages [61, 10, 15, 9] and unsupervised image translation [35, 68]. Our vec2vec method uses adversarial losses and cycle consistency to learn to encode embeddings into a shared latent space and decode with minimal loss. This makes unsupervised translation possible. We use a basic adversarial approach with vector space preservation [45] to learn a mapping from an unknown embedding distribution to a known one.
vec2vec is the first method to successfully translate embeddings from the space of one model to another without paired data. [^1] Prior work has successfully translated word embeddings between languages, typically relying on overlapping vocabularies across languages. In contrast, we translate embeddings of entire sequences between model spaces. vec2vec translations achieve cosine similarity as high as $0.92$ to the ground-truth vectors in their target embedding spaces and perfect matching on over $8000$ shuffled embeddings (without access to the set of possible matches in advance).
@@ 2 Problem formulation: unsupervised embedding translation
We do assume access to a different encoder $M_{2}$ that we can query at will to generate new embeddings in some other space. We also assume high-level distributional knowledge about the hidden documents: their modality (text) and language (e.g., English). To extract information, we may translate $\{u_{1},\ldots u_{n}\}$ into the output space of $M_{2}$ and apply techniques like inversion that take advantage of the encoder.
Limitations of correspondence methods. There is significant prior research on the problem of matching or correspondence between sets of embedding vectors [1, 48, 8, 53]. These methods typically assume that the two (or more) sets of embeddings are generated by different encoders on the same or highly-overlapping inputs. In other words, for each unknown vector, there must already exist a set of candidate vectors in a different embedding. In practice, it is unrealistic to expect that such a database be available, so these methods are not directly applicable. Some matching methods, however, support translation between embedding spaces without overlapping inputs. Our experiments demonstrate that these methods struggle significantly, even when correspondence exists.
@@
The Strong Platonic Representation Hypothesis. Our hope that unsupervised embedding translation is possible at all rests on the stronger version of the Platonic Representation Hypothesis [18]. Our conjecture is as follows: neural networks trained with the same objective and modality, but with different data and model architectures, converge to a universal latent space such that a translation between their respective representations can learned without any pairwise correspondence.
Translation enables information extraction. Solving unsupervised translation will allow us to use information extraction tools designed to operate on vectors produced by known encoders. For example, we could apply inversion models [42, 66] to recover unknown documents $\{d_{i}\}$.
@@ 3 Our method: vec2vec
Unsupervised translation has been successful in computer vision, using a combination of cycle consistency and adversarial regularization [35, 68]. Our design of vec2vec is inspired in part by these methods. We aim to learn embedding-space translations that are cycle-consistent (mapping to and from an embedding space should end in the same place) and indistinguishable (embeddings for the same text from either space should have identical latents).
@@ 3.2 Optimization
Adversarial loss. The adversarial loss encourages generated embeddings to match the empirical distributions of original embeddings both at the embedding and latent levels. Specifically, applying the standard GAN loss formulation [13] to both levels yields:
@@
Generator. Because adversarial losses alone do not guarantee that translated embeddings preserve semantics [68], we introduce three additional constraints to help the generator learn a useful mapping:
@@
Vector space preservation (VSP) ensures that pairwise relationships between generated embeddings remain consistent under translation [45, 64]. Given a batch of $B$ embeddings $x_{1},...,x_{B}$ and $y_{1},...,y_{B}$, we sum their average pairwise distances after translation by both $F_{1}$ and $F_{2}$:
$\mathcal{L}_{\text{VSP}}(F_{1},F_{2})=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}\bigg[$ $\|M_{1}(x_{i})\cdot M_{1}(x_{j})-F_{1}(M_{1}(x_{i}))\cdot F_{1}(M_{1}(x_{j}))\|_{2}^{2}$
$+\|M_{2}(y_{i})\cdot M_{2}(y_{j})-F_{2}(M_{2}(y_{i}))\cdot F_{2}(M_{2}(y_{j}))\|_{2}^{2}\bigg]$
@@ 4.1 Preliminaries
Datasets. We use the Natural Questions (NQ) [25] dataset of user queries and Wikipedia-sourced answers for training (a $2$-million subset) and evaluation (a $65536$ subset). To evaluate information extraction, we use TweetTopic [2], a dataset of tweets multi-labeled by 19 topics; a random $8192$-record subset of Pseudo Re-identified MIMIC-III (MIMIC) [28], a pseudo re-identified version of the MIMIC dataset [19] of patient records multi-labeled by 2673 MedCAT [24] disease descriptions; and a random $50$-email subset of the Enron Email Corpus (Enron) [21], an unlabeled, public dataset of internal emails of a defunct energy company.
Models. Table˜1 lists the embedding models representing three size categories, four transformer backbones, and two output dimensionalities. Granite is multilingual, CLIP is multimodal.
@@
[46] gtr 110 T5 2021 768 512
[49] clip 151 CLIP 2021 512 77
[57] e5 109 BERT 2022 768 512
[32] gte 109 BERT 2023 768 512
[67] stella 109 BERT 2023 768 512
[14] granite 278 RoBERTa 2024 768 512
@@
Training. Unless otherwise specified, each vec2vec is trained on two sets of embeddings generated from disjoint sets of $1$ million $64$-token sequences sampled from NQ (see Appendix˜D for experiments with fewer embeddings). Due to GAN instability [52], we select the best of multiple initializations and leave more robust training to future work.
@@ 4.2 Evaluating translation
vec2vec is the first unsupervised embedding translator, thus there is no direct baseline. As our Naïve baseline, we simply use $F(x)=x$ to measure geometric similarity between embedding spaces. The second (pseudo)baseline is Oracle-aided optimal transport. It assumes that candidate targets are known and is thus strictly easier than vec2vec and the Naïve baseline. We solve optimal assignment, $\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{n}\cos(u_{i},v_{\pi(i)})$, via either the Hungarian, Earth Mover’s Distance, Sinkhorn, or Gromov-Wasserstein algorithms, choosing the solver with the lowest rank for each experiment.
@@ 4.3 Evaluating information extraction
The second method is embedding inversion that recovers text inputs from embeddings. Since [42] requires a pre-trained inversion model for each embedding space, we use [66] instead to generate an approximation $d^{\prime}_{i}$ of $d_{i}$ from $F(M_{1}(d_{i}))$ in a zero-shot manner. We measure the extracted information using LLM judge accuracy: the fraction of translated embeddings for which GPT-4o determines that $d^{\prime}$ reveals information in $d$. See Appendix˜E for our prompt.
@@
$M_{1}$ $M_{2}$ $\cos(\cdot)$ $\uparrow$ Top-1 $\uparrow$ Rank $\downarrow$ $\cos(\cdot)$ $\uparrow$ Top-1 $\uparrow$ Rank $\downarrow$
@@
Table˜3 shows that this performance extends to out-of-distribution data. Our vec2vec translators were trained on NQ (drawn from Wikipedia), yet exhibit high cosine similarity, high accuracy, and low rank when evaluated on tweets (which are far more colloquial and use emojis) and medical records (which contain domain-specific jargon unlikely to appear in NQ). In Appendix˜C, we show that baseline methods fail on cross-backbone embedding pairs.
@@
Finally, Table˜4 shows that vec2vec can even translate to and from the space of CLIP, a multimodal embedding model which was trained in part on image data. While the translations are not as strong as in Table˜2, vec2vec consistently outperforms the optimal transport baseline. These results show the promise of our method at adapting to new modalities: in particular, the embedding space of CLIP has been successfully connected to other modalities such as heatmaps, audio, and depth charts [12].
@@ 6 Using vec2vec translations to extract information
Zero-shot inversion. Inversion, i.e., reconstruction of text inputs, is more ambitious than attribute inference. vec2vec translations retain enough semantic information that off-the-shelf, zero-shot inversion methods like [66], developed for embeddings computed by standard encoders, extract information for as many as 80% of documents given only their translated embeddings, for some model pairs (Figure˜5). These inversions are imperfect and we leave development of specialized inverters for translated embeddings to future work. Nevertheless, as exemplified in Figure˜6, they still extract information such as individual and company names, dates, promotions, financial information, outages, and even lunch orders. In Appendix˜E, we show the prompt we use to measure extraction.
@@
Figure 6: Examples of inversions that infer entities and content.
@@ 7 Related work
Representation alignment. Similarities between representations of different neural networks are investigated in [26, 31, 58, 5, 18, 60, 30]. Methods based on CCA [41], SVCCA, [50], CKA [23, 37], ICA [62], time-series [38], and GUIs [16] have been used to compare embeddings from different subspaces. [36, 44, 39, 56, 47] harness representation similarity for zero-shot stitching, substitution, domain transfer, and multi-modal adaptation. All rely on some amount of paired data, which is difficult to reduce [6]. Our method does not just measure similarity, we learn how to translate representations across spaces without any paired data.
Optimal transport. The problem of unsupervised optimal transport has been studied for image style transfer [17, 35, 68], word translation [61, 10, 15, 9, 20], and natural language sequence translation [51, 27, 1, 4, 63, 3]. Our method builds on these works, which often employ a combination of cycle-consistency and adversarial loss. Importantly, unlike prior word and sequence translation methods, multiple representations of the same input (e.g., heavily overlapping word vocabularies) are unavailable in our setting. [53] proposes a solver for matching small sets of embeddings between different vision-language models. Our method goes well beyond matching by taking unknown embeddings and generating matching embeddings in the space of another model.
Embedding inversion. An emerging line of research investigates decoding text from language model embeddings [54, 29, 42] and outputs [43, 7, 65]. vec2vec helps apply these to unknown embeddings, without an encoder or paired data, by translating them to the space of a known model.
Bridging modality gaps. Previous work has noted an inherent “gap” between image- and text-based models [33] and proposed various ways to unify the modalities [55]. Some approaches feed image embeddings directly into language models [22, 59, 11, 34], while others generate captions from image embeddings [40] or even from text embeddings themselves [42]. [12] introduces a shared embedding space that integrates inputs from multiple modalities, including text, audio, and vision. In contrast, our post-hoc approach directly translates between representations and complements these systems by enabling inputs from a wide variety of embedding models.
@@ 8 Discussion and Future Work
This research is supported in part by the Google Cyber NYC Institutional Research Program. JM is supported by the National Science Foundation.
@@ Appendix A Compute
For our main experiments, we trained 36 distinct vec2vec models, with training durations ranging from 1 to 7 days per model, depending on the specific GPU and model pair (which affected convergence rates). Additional 9 vec2vec models were trained for our ablations, bringing the total to 45 vec2vec models. Taking a conservative estimate of the average training time, this amounted to approximately 180 GPU days (45 models $\times$ 4 days / model).
@@
  • For each of our 45 vec2vec models, our full evaluation on NQ, TweetTopic, and MIMIC datasets required approximately 1 hour per model across all GPU types.
  • The 36 non-ablation GANs underwent additional analysis on the Enron email corpus, requiring about 1.5 hours for inversion and downstream LLM evaluation.
  • Naive baselines (30 models without CLIP) required approximately 30 minutes each for evaluation, roughly half the time needed for our full models.
  • Optimal transport baselines (36 models and 4 different algorithms) were run exclusively on CPU, each requiring approximately 4 hours of computation time.
In total, our experiments consumed approximately 180 GPU days for training and an additional 78 GPU hours for evaluation and analysis. There were approximately 144 CPU hours required for optimal transport.
Appendix B Extended oracle-aided optimal transport baseline results

### v4

@@ Abstract
Figure 1: Left: input embeddings from different model families (T5-based GTR [undefat] and BERT-based GTE [undefae]) are fundamentally incomparable. Right: given unpaired embedding samples from different models on different texts, our model learns a latent representation where they are closely aligned.
@@ 1 Introduction
The Platonic Representation Hypothesis [undefq] conjectures that all vision models of sufficient size converge to the same latent representation. We propose a stronger, constructive version of this hypothesis for text models: the universal latent structure of text representations can be learned and, furthermore, harnessed to translate representations from one space to another without any paired data or encoders.
@@
We draw inspiration from research on aligning word embeddings across languages [undefaai, undefi, undefn, undefh] and unsupervised image translation [undefai, undefaaq]. Our vec2vec method uses adversarial losses and cycle consistency to learn to encode embeddings into a shared latent space and decode with minimal loss. This makes unsupervised translation possible. We use a basic adversarial approach with vector space preservation [undefas] to learn a mapping from an unknown embedding distribution to a known one.
vec2vec is the first method to successfully translate embeddings from the space of one model to another without paired data. [^1] Prior work has successfully translated word embeddings between languages, typically relying on overlapping vocabularies across languages. In contrast, we translate embeddings of entire sequences between model spaces. vec2vec translations achieve cosine similarity as high as $0.96$ to the ground-truth vectors in their target embedding spaces and perfect matching on over $8000$ shuffled embeddings (without access to the set of possible matches in advance).
@@ 2 Problem formulation: unsupervised embedding translation
We do assume access to a different encoder $M_{2}$ that we can query at will to generate new embeddings in some other space. We also assume high-level distributional knowledge about the hidden documents: their modality (text) and language (e.g., English). To extract information, we may translate $\{u_{1},\ldots u_{n}\}$ into the output space of $M_{2}$ and apply techniques like inversion that require the encoder.
Limitations of correspondence methods. There is significant prior research on the problem of matching or correspondence between sets of embedding vectors [undef, undefav, undefg, undefaaa]. These methods typically assume that the two (or more) sets of embeddings are generated by different encoders on the same or highly-overlapping inputs. In other words, for each unknown vector, there must already exist a set of candidate vectors in a different embedding. In practice, it is unrealistic to expect that such a database be available, so these methods are not directly applicable. Some matching methods, however, support translation between embedding spaces without overlapping inputs. Our experiments demonstrate that these methods struggle significantly, even when correspondence exists.
@@
The Strong Platonic Representation Hypothesis. Our hope that unsupervised embedding translation is possible at all rests on the stronger version of the Platonic Representation Hypothesis [undefq]. Our conjecture is as follows: neural networks trained with the same objective and modality, but with different data and model architectures, converge to a universal latent space such that a translation between their respective representations can be learned without any pairwise correspondence.
Translation enables information extraction. Solving unsupervised translation will allow us to use information extraction tools designed to operate on vectors produced by known encoders. For example, we could apply inversion models [undefap, undefaan] to recover unknown documents $\{d_{i}\}$.
@@ 3 Our method: vec2vec
Unsupervised translation has been successful in computer vision, using a combination of cycle consistency and adversarial regularization [undefai, undefaaq]. Our design of vec2vec is inspired in part by these methods. We aim to learn embedding-space translations that are cycle-consistent (mapping to and from an embedding space should end in the same place) and indistinguishable (embeddings for the same text from either space should have identical latents).
@@ 3.2 Optimization
Adversarial. The adversarial loss encourages generated embeddings to match the empirical distributions of original embeddings both at the embedding and latent levels. Specifically, applying the standard GAN loss formulation [undefl] to both levels yields:
@@
Generator. Because adversarial losses alone do not guarantee that translated embeddings preserve semantics [undefaaq], we introduce three additional constraints to help the generator learn a useful mapping:
@@
Vector space preservation (VSP) ensures that pairwise relationships between translated embeddings are consistent with the target space [undefas, undefaal]. Given a batch of $B$ embeddings $x_{1},...,x_{B}$ and $y_{1},...,y_{B}$, we sum their average pairwise distances after translation by both $F_{1}$ and $F_{2}$:
$\mathcal{L}_{\text{VSP}}(F_{1},F_{2})=\frac{1}{B^{2}}\sum_{i=1}^{B}\sum_{j=1}^{B}\bigg[$ $\|M_{1}(x_{i})\cdot M_{1}(x_{j})-F_{2}(M_{2}(y_{i}))\cdot F_{2}(M_{2}(y_{j}))\|_{2}^{2}$
$+\|M_{2}(y_{i})\cdot M_{2}(y_{j})-F_{1}(M_{1}(x_{i}))\cdot F_{1}(M_{1}(x_{j}))\|_{2}^{2}\bigg]$
@@ 4.1 Preliminaries
Datasets. We use the Natural Questions (NQ) [undefx] dataset of user queries and Wikipedia-sourced answers for training (a $2$-million subset) and evaluation (a $65536$ subset). To evaluate information extraction, we use TweetTopic [undefa], a dataset of tweets multi-labeled by 19 topics; a random $8192$-record subset of Pseudo Re-identified MIMIC-III (MIMIC) [undefaa], a pseudo re-identified version of the MIMIC dataset [undefr] of patient records multi-labeled by 2673 MedCAT [undefw] disease descriptions; and a random $50$-email subset of the Enron Email Corpus (Enron) [undeft], an unlabeled, public dataset of internal emails from a defunct energy company. In Appendix˜D, we ablate a model on MS COCO [undefag], a captioned image dataset, to evaluate performance on multimodal retrieval.
Models. Table˜1 lists the embedding models representing four size categories, five transformer backbones, and two output dimensionalities. Granite is multilingual; CLIP is multimodal. Since Qwen is very compute-intensive, we only evaluate it for a single model pair in Appendix˜C.
@@
[undefat] gtr 110 T5 2021 768 512
[undefaw] clip 151 CLIP 2021 512 77
[undefaae] e5 109 BERT 2022 768 512
[undefae] gte 109 BERT 2023 768 512
[undefaao] stella 109 BERT 2023 768 512
[undefm] granite 278 RoBERTa 2024 768 512
[undefaap] qwen 4000 Qwen3 2025 2560 32K
@@
Training. Unless otherwise specified, each vec2vec is trained on two sets of embeddings generated from disjoint sets of $1$ million $64$-token sequences sampled from NQ (see Section˜7 for experiments with fewer embeddings). Due to GAN instability [undefaz], we select the best of multiple initializations (see Appendix˜E) and leave more robust training to future work. See Appendix˜A for compute details.
@@ 4.2 Evaluating translation
vec2vec is the first unsupervised embedding translator, thus there is no direct baseline. As our Naïve baseline, we simply use $F(x)=x$ to measure geometric similarity between embedding spaces. The second (pseudo)baseline is Oracle-aided optimal transport. It assumes that candidate targets are known and is thus strictly easier than vec2vec and the Naïve baseline. We solve optimal assignment, $\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{n}\cos(u_{i},v_{\pi(i)})$, via either the Hungarian, Earth Mover’s Distance, Sinkhorn, or (Entropic) Gromov-Wasserstein algorithms, choosing the solver with the lowest rank for each experiment. See Appendix˜B for more details.
@@ 4.3 Evaluating information extraction
The second method is embedding inversion that recovers text inputs from embeddings. Since [undefap] requires a pre-trained inversion model for each embedding space, we use [undefaan] instead to generate an approximation $d^{\prime}_{i}$ of $d_{i}$ from $F(M_{1}(d_{i}))$ in a zero-shot manner. We measure the extracted information using LLM judge accuracy: the fraction of translated embeddings for which GPT-4o determines that $d^{\prime}$ reveals information in $d$. See Appendix˜H for our prompt.
@@
Table˜3 shows that this performance extends to out-of-distribution data. Our vec2vec translators were trained on NQ (drawn from Wikipedia), yet exhibit high cosine similarity, high accuracy, and low rank when evaluated on tweets (which are far more colloquial and use emojis) and medical records (which contain domain-specific jargon unlikely to appear in NQ). In Appendix˜F, we show that baseline methods fail on cross-backbone embedding pairs.
@@
Finally, Table˜4 shows that vec2vec can even translate to and from the space of CLIP, a multimodal embedding model which was trained in part on image data. While the translations are not as strong as in Table˜2, vec2vec consistently outperforms the optimal transport baseline. These results show the promise of our method at adapting to new modalities: in particular, the embedding space of CLIP has been successfully connected to other modalities such as heatmaps, audio, and depth charts [undefk].
@@ 6 Using vec2vec translations to extract information
Zero-shot inversion. Inversion, i.e., reconstruction of text inputs, is more ambitious than attribute inference. vec2vec translations retain enough semantic information that off-the-shelf, zero-shot inversion methods like [undefaan], developed for embeddings computed by standard encoders, extract information for as many as 80% of emails and 67% of tweets given only their translated embeddings, for some model pairs (Figure˜5 and Appendix˜G). These inversions are imperfect and we leave development of specialized inverters for translated embeddings to future work. Nevertheless, as exemplified in Figure˜6, they still extract potentially sensitive information such as individual and company names, dates, promotions, financial information, outages, and even lunch orders. In Appendix˜H, we show the prompt we use to measure extraction.
@@
Figure 6: Examples of Enron Email Corpus inversions that infer entities and content.
7 Ablations
Method $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$
vec2vec 0.75 (0.0) 0.91 2.64 (0.1)
Naïve Baseline 0.04 (0.0) 0.00 4084.15 (9.2)
OT Baseline 0.70 (0.0) 0.00 3064.16 (8.9)
– VSP loss 0.58 (0.0) 0.00 4196.64 (9.2)
– CC loss 0.50 (0.0) 0.00 3941.36 (9.3)
– latent GAN 0.49 (0.0) 0.00 3897.09 (9.5)
– VSP and CC loss 0.47 (0.0) 0.00 3365.24 (9.3)
– hyperparam. tuning 0.50 (0.0) 0.00 4011.73 (9.3)
Each component of our method is important. We ablate our method subtractively, measuring the key metrics after removing individual components of our algorithm (described in Section˜3). Table˜6 shows that each component appears to be critical to building good translations. While vec2vec’s $\cos(\cdot)$ is higher than the naïve baseline, it performs worse across the board than the OT baseline and does not preserve the geometry of the vector space.
NN $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$
1000000 0.75 (0.0) 0.92 2.73 (0.2)
10000 0.57 (0.0) 0.01 1462.21 (20.)
50000 0.74 (0.0) 0.81 3.91 (0.6)
100000 0.74 (0.0) 0.85 4.52 (0.4)
500000 0.75 (0.0) 0.92 2.73 (0.2)
vec2vecs can be trained with significantly less data. In Sections˜5 and 6, we use 1M-point subsets of NQ to train our vec2vec models. Now, we train the gte $\to$ gtr vec2vec with 1M GTR embeddings but fewer GTE embeddings. Table˜7 shows that with as few as 10K embeddings, the translators still learn something (i.e. are better than random). Translations trained on 50K embeddings are almost as good as those trained on 1M. Translations generally improve with more training data.
@@ 8 Related work
Representation alignment. Similarities between representations of different neural networks are investigated in [undefy, undefad, undefaaf, undefd, undefq, undefaah, undefac]. Methods based on CCA [undefao], SVCCA, [undefax], CKA [undefv, undefak], ICA [undefaaj], time-series [undefal], and GUIs [undefo] have been used to compare embeddings from different subspaces. [undefaj, undefar, undefam, undefaad, undefau] harness representation similarity for zero-shot stitching, substitution, domain transfer, and multimodal adaptation. All rely on some amount of paired data, which is difficult to reduce [undefe]. Our method does not just measure similarity, we learn how to translate representations across spaces without any paired data.
Optimal transport. The problem of unsupervised optimal transport has been studied for image style transfer [undefp, undefai, undefaaq], word translation [undefaai, undefi, undefn, undefh, undefs], and natural language sequence translation [undefay, undefz, undef, undefc, undefaak, undefb]. Our method builds on these works, which often employ a combination of cycle-consistency and adversarial loss. Importantly, unlike prior word and sequence translation methods, multiple representations of the same input (e.g., heavily overlapping word vocabularies) are unavailable in our setting. [undefaaa] proposes a solver for matching small sets of embeddings between different vision-language models. Our method goes well beyond matching by taking unknown embeddings and generating matching embeddings in the space of another model.
Embedding inversion. An emerging line of research investigates decoding text from language model embeddings [undefaab, undefab, undefap] and outputs [undefaq, undeff, undefaam]. vec2vec helps apply these to unknown embeddings, without an encoder or paired data, by translating them to the space of a known model.
Bridging modality gaps. Previous work has noted an inherent “gap” between image- and text-based models [undefaf] and proposed various ways to unify the modalities [undefaac]. Some approaches feed image embeddings directly into language models [undefu, undefaag, undefj, undefah], while others generate captions from image embeddings [undefan] or even from text embeddings themselves [undefap]. [undefk] introduces a shared embedding space that integrates inputs from multiple modalities, including text, audio, and vision. In contrast, our post-hoc approach directly translates between representations and complements these systems by enabling inputs from a wide variety of embedding models.
@@ 9 Discussion and Future Work
This research is supported in part by the Google Cyber NYC Institutional Research Program. RJ is supported by the Digital Life Initiative Fellowship and JM by the National Science Foundation.
@@ Appendix A Compute
For our final results, we trained 25 vec2vec models fully and 30 models partially (see Appendix˜E). The full models’ training durations usually ranged from 1 to 7 days, depending on the specific GPU and model pair (which affected convergence rates). Partial convergence was stopped after 2 days. Due to the size of Qwen, our (qwen, gte) ablation was trained for 20 days on an A100. Taking a conservative estimate of the average training time, this amounted to approximately 176 GPU days (24 models $\times$ 4 days / model + 30 models $\times$ 2 days / model + 1 (qwen, gte) $\times$ 20 days / model).
@@
  • The 10 main vec2vec models required $\sim$1 hour each for NQ, TweetTopic, and MIMIC evaluation (across GPU types), plus 30 minutes for attribute extraction on TweetTopic and MIMIC, and 1.5 hours for inversion and downstream LLM evaluation on Enron and TweetTopic. Naive baselines required $\sim$30 minutes each across all datasets.
  • The 15 additional fully-trained models required 30 minutes each for NQ evaluation, with an extra 30 minutes for MS COCO evaluation of (clip, granite).
  • Optimal transport baselines ran on CPU only, requiring $\sim$1 hour per dataset (three datasets for main models, one for others).
In total, our experiments consumed almost 176 GPU days for training and an additional 42 GPU hours for evaluation and analysis. An additional 45 CPU hours were required for optimal transport.
Appendix B Oracle-aided optimal transport baseline